### v1

@@ Abstract
This paper investigates the nuanced interplay of gender, birth order, and innate ability in shaping educational disparities among children within households, employing both a reduced-form analysis and a structural model of household resource allocation. By decomposing overall disparities, I identify the contributions of gender, birth order, and innate ability differences. In the context of Benin, for households with non-educated heads and mixed-gender children, total inequality comprises 50% gender disparity, 20% birth order effect, and 30% ability gap. Conversely, in households with only sons or only daughters and non-educated heads, total inequality is predominantly driven by ability disparities—70% for daughters and 83% for sons, with lesser contributions from birth order effects. Furthermore, my analysis reveals that in households with non-educated heads, firstborn daughters surpass their younger brothers in educational attainment on the intensive margin if their innate ability exceeds their brothers’ by at least 13%, a figure reduced to 8% in households with college-educated parents. Additionally, the study unveils parental preferences favoring sons’ education over daughters’ among non-educated parents, with a perceived 22% higher average benefit. Targeted policies aimed at reducing composite education costs prove effective in mitigating gender and birth order gaps, albeit with a modest 5% reduction in total inequality. Overall, this research underscores the complex dynamics influencing intra-household educational inequalities and suggests policy avenues to address them.
@@ 1 Introduction
The trade-off between the number of children one chooses to have and the quality of those children has been a subject of extensive research. This exploration began with the seminal contributions of De Tray (1970) and Becker and Lewis (1973). The quantity-quality trade-off model delves into how changes in the quantity of children can impact their average quality. Numerous empirical studies have provided support for this model in various contexts, consistently revealing a negative relationship between the number of children, often used as an empirical measure of quantity, and the educational attainment of those children, typically used as an empirical measure of quality. These findings align with the predictions put forth by the theoretical framework presented by Becker and Lewis (1973) and have been substantiated by research such as Montgomery (1995) and Li et al. (2008). The current state of the literature on fertility choices (Conley and Glauber (2006), Maralani (2008), Li et al. (2008), Weng et al. (2019)) states that rich and educated families tend to have fewer children while allocating greater investments in the education of those children, in comparison to less affluent and less educated families. In other words, the average level of education among children varies across different families, influenced by endogenous factors such as family size and exogenous factors like the socio-economic background of the parents.
@@
When we extend the analysis beyond the choice of the aggregate quality of children to how it is distributed among children, we observe that, not only does the average change across households, but so does the variance of quality. In addition, the variance is most likely non-zero for a majority of households when there are budget constraints and no compulsory education laws. This presumption of a non-zero variance is justified by evidence of the influence of factors such as gender, birth order and innate ability on education attainment of children. The educational outcomes of children are strongly influenced by their individual characteristics, leading to disparities in the amount of education they receive. Several studies have shed light on these disparities, with findings indicating that girls tend to receive less education due to factors such as gender bias or gender preference (as documented by Biswas (2000) and Ota and Moffatt (2007)). Additionally, birth order can play a role in the educational opportunities afforded to children, with elder siblings benefiting or facing disadvantages, as observed in studies like Ota and Moffatt (2007), Weng et al. (2019), Fergusson et al. (2006), De Haan (2010), Moshoeshoe et al. (2016), and Esposito et al. (2020). Furthermore, children with higher abilities are more likely to receive increased educational opportunities and education attainment, as suggested by the research of Becker and Tomes (1976), Dizon-Ross (2019), and Giannola (2023). However, there exists a notable gap in the research landscape, as there is a scarcity of studies that comprehensively analyze all these various sources of disparities within the same analytical framework. Such a framework, capable of simultaneously examining gender-based differences, birth order effects, and the impact of individual abilities on educational inequality within households, holds the potential for interesting counterfactual and education policies analyses. Specifically, it enables us to explore how these factors interact with one another to shape inequality. Furthermore, this comprehensive approach is crucial for estimating the proportion of inequality attributable to disadvantages and the portion caused by ability-based resource allocation within households. The former allows us to design education policies which target inequality due gender or birth order disadvantages.
In this paper, I use reduced form approach and suggest a structural model to break down the overall intra-household differences in children’s education into parts related to gender, birth order, and ability. Firstly, the model helps us find out if the total disparities observed is solely due to biases from parents and society or because there is high variation in children’s abilities. Secondly, it allows us to measure how much more intelligent a daughter (or firstborn) needs to be in order to have the same educational opportunities as a son (or second born). Lastly, using this model, we can examine how differences in education of children within a family change with a cost reduction policy compared to a policy which aim to change parents’ preference.
I estimate the model using data from Benin in West Africa. Benin is a good place to apply this model for three main reasons. First, even though there is a law stipulating that everybody must finish at least primary school, it’s not always followed by parents or enforced by the government. Second, there are big differences in how much education individuals in the same household get. Also, these differences can be very different from one family to another. Lastly, there is evidence that girls and firstborns often don’t get the same education opportunities as others. For this exercise, I focused on households where there were only two adult children still residing with their parents. Among those households there is difference in how much education resources they have to distribute and the education attainment of the head of household. Taking that into account, I applied my method to households with different characteristics separately.
My analysis uncovered three important findings. First, when it comes to educating daughters compared to their brothers within a household with a non-educated head of household, large part (70%) of the differences in the education of brothers and sisters is due to disadvantages rather than variations in abilities. That number is 33% among college educated parents. However, when examining the education of firstborns compared to their siblings of the same gender, more than half (70% for female and 80% for male) of the disparities are attributed to differences in abilities. These results suggests that in some households gender disadvantage reduces disparities in children’s education through the reduction of the extent to which the talent of a child affect their education. This implies that removing/reducing gender and birth order disadvantages will not systematically reduce inequality for all households. It will increase the amount of disparities in some households.
Second, When we reduce education costs for firstborn children and daughters, we can decrease the differences in education within certain families. This is especially true in families where the firstborn children and daughters have lower abilities compared to their siblings or where parents don’t really take the children’s abilities into consideration when making decisions about education. However, if the firstborn children and daughters are actually more capable than their siblings and parents do consider their abilities, a policy that eases the financial burden of schooling for them might unintentionally increase educational differences within the family. This brings up a dilemma about finding the right balance between giving everyone equal educational opportunities, regardless of factors like gender, birth order, and ability, and ensuring that children with similar abilities get the same opportunities, regardless of gender and birth order. In the latter case we are allowing for rewarding high abilities but not for penalizing gender and birth order.
In terms of policy changes, it seems more implementable to aim for the latter option. That means focusing on policies that increase the net return from education for firstborn children and daughters either through schooling cost reduction or reduction of salary gap between male and female. These policies will help decrease the impact of gender and birth order on educational differences. On the other hand, it’s trickier to implement a policy that reduces how much parents consider a child’s abilities in distributing the few resources that they have, when facing financial challenges. Especially given that it is more costly to give high level formal education to a low ability child, and that there are other more suitable alternatives such as apprenticeship.
Third, When all kids in a family are considered equally capable, the differences in education between them created by gender and birth order is at its highest point. This happens because parents see the benefit from their investment in education and the child’s ability as complementary. In other words, parents think that spending more on the education of a child with higher ability will bring greater benefits compared to investing in a child with lower ability. So, if the firstborn or a daughter has lower abilities than their siblings, they will receive less investment in education even without considering gender and birth order disadvantages. This means that the disadvantages they face because of gender and/or birth order have a smaller impact on increasing the overall disparities in education in term of magnitude. On the flip side, if the firstborn or a daughter is more capable than their siblings, they would get more investment in their education compared to their siblings, in the absence of gender and birth order disadvantages. As a result, the disadvantage they face due to gender and birth order leads to a decrease in the educational differences within the family. This decrease becomes less pronounced as their intelligence level increases.
@@
The subsequent sections of this paper are structured as follows. In Section 2, I present the literature review along with the specific contributions of this paper. Section 3 offers an overview of the data utilized for the estimation of the model, while Section 4 presents key empirical observations and facts derived from the analysis. Section 5 is dedicated for the model’s setup, outlining the identification strategy for key parameters, and describing the inference and estimation procedures employed in this study. Lastly, in Section 6, I present counterfactual analysis to further explore the implications of my findings. Section 7 presents a robustness analysis of the key parameters of interest, and Section 8 concludes the paper.
2 Literature Review
This paper contributes to the extensive literature on fertility choice models and the quantity-quality trade-off faced by households within a country. Early developed theoretical models (DeTray (1970), Becker and Lewis (1973)) predicts that quantity and quality of children are negatively related due to the observation that the shadow price of quantity is influenced by quality, and vice versa. More specially, they have shown that a reduction in the quantity of children results in a decrease in the shadow price of quality. Consequently, this decline in the shadow price of quality leads to an increase in the overall quality of the children. The current state of the literature on fertility choices (Montgomery (1995), Conley and Glauber (2006), Maralani (2008), Li et al. (2008), Weng et al. (2019)) states that rich and educated families tend to have fewer children while allocating greater investments in the education of those children, in comparison to less affluent and less educated families. This trend reflects a significant pattern where socio-economic factors play a crucial role in shaping fertility decisions and educational investments within households. However, some studies still found a positive or no correlation between fertility and education of children either for rural or old households (Montgomery (1995), Black et al. (2005), Maralani (2008)). This is often explained by the lack of contraception methods and poor public education systems. This paper contributes to that strand of the literature by offering two key contributions. Firstly, it empirically demonstrates the existence of a quantity-quality trade-off within the context of Benin. Second, it builds on this result to relax the assumption of equal education for children in the same household.
@@ 2 Literature Review
This paper also contributes to the literature on the within-household schooling decision, particularly factors influencing parents’ distribution decision of education resources among siblings. One key determinant of these distribution decisions is the gender or gender composition of households. Previous research has shown that daughters are less likely to receive education or have lower educational attainment on average. Studies have shown that, while the presence of elder sisters tends to increase the likelihood of schooling, the presence of younger brothers may decrease it (Biswas (2000), Ota and Moffatt (2007), Ombati and Ombati (2012), Osadan and Burrage (2014), Psaki et al. (2018)). Another influential factor is the birth order of children, with mixed findings in previous studies. Some papers suggest a positive effect of birth order on children’s education (Ota and Moffatt (2007), Weng et al. (2019)), while others have shown that later-born children have lower educational attainment (Fergusson et al. (2006), De Haan (2010), Moshoeshoe et al. (2016), Esposito et al. (2020)). Finally, a child’s innate ability or talent plays a role in parental distribution decisions. Studies have demonstrated that parents invest more in the human capital of endowed children and allocate more nonhuman capital to less endowed children (Becker and Tomes (1976), Dizon-Ross (2019), Giannola (2023)). When parents are compelled to invest in the nonhuman capital of low-ability children, this leads to an inefficient equilibrium, where the investment in the human capital of high-ability children is not optimized (Nerlove et al. (1984)). This paper adds to this existing literature in two significant ways. Firstly, it examines a context where parents are not constrained to compensate lower ability children by investing in their nonhuman capital but, instead, rely on family taxes (Wantchekon et al. (2015)). Secondly, this paper develops a household resources distribution structural model that allows for a more flexible analysis of the distribution of education resources within the household. In this model, the assumption of equal distribution is relaxed, enabling a detailed exploration of the interactions between gender and, birth order disadvantages, and the innate abilities of individual children in influencing parental distribution decisions.
@@
Finally, this paper contributes to the literature on within household inequality in children’s human capital (Giannola (2023)). Giannola (2023) has shown in the context of India that observed inequality within households is partly explained by parents investing more in the human capital of high-achieving children, especially when they are financially constrained. This behavior stems from the fact that parents are not particularly averse to inequality and tend to reinforce the gap in learning created by innate ability rather than correct it. This paper contributes to that literature by first building upon the result that parents unequally invest in the human capital of high-achieving children in contexts where the education system is better tailored to serve high-achieving students. Second, this paper interacts with that result and examines how it relates to other sources of unequal distribution, such as gender and birth order.
@@ 3 Data Description and Definition of Key Variables
In this section, I present the data used to analyze the question of interest in this paper. The context of Benin is particularly compelling for analyzing the interaction between gender disadvantage, birth order advantage, and innate ability in shaping within household inequality for three reasons. First, there are a lot of disparities in education attainment within households due to the lack of compulsory education law coupled with financial constraints and the availability of outside options other than formal education such as child labor within the households and apprenticeship. Second, there is factual evidence of gender and birth order disadvantages. Lastly, there is observational evidence that despite the prevalence of gender disadvantages in the population, daughters in some households get higher education than their brothers.
@@ 3.1 Sample and Data
The 2013 Population and Habitation Census data of Benin; used in this paper; offers a comprehensive insights into households and their members living in the country during that year. Conducted by the National Bureau of Statistics of Benin, this census provides data at both the household and individual levels. For the purpose of this paper, the focus is directed towards individuals who identify themselves as the children of household heads, enabling to get information on parental  [^3] Parents here refers to one of the parents, either the mother or the father. This because it is not possible to have both for household with single parents and to identify the biological mother for polygamous households. and sibling characteristics. To be specific, the variable ”Number of children” represents the observed number of children within each household  [^4] It does not include children who moved out of the family house before the census.. For the primary analysis, only households with children aged between 25 and 40 years are included. This age range is chosen to ensure that the children have either completed their education or nearly achieved complete educational attainment.
The inequality analysis focuses on households with at least 2 such children falling within the specified age range and at least one child with some educational attainment. This specific sample selection is motivated by the desire to examine the consequences of providing equal education to all children, as opposed to the alternative of not educating any children. The resulting sample comprises approximately $90,000$ individuals and $32,000$ households, serving as the basis for further investigation.
@@ 3.2 Key variables and measurement
The data set contains several key variables used in this paper, including gender, age, religion, area of residence, family size, household wealth index, and educational attainment of individuals, as well as their parents’ and a subset of their siblings’ variables. Apart from these variables, I also created measures for within household inequality, within household average years of education, and gender composition of children within a household. A description of each variable and their measurement is as follows:
Inequality: It refers to inequality in the education attainment of children within a given household. It is measured by the within household standard deviation and range of children’s education attainment. These two measures are used interchangeably throughout this paper.
High school graduate head of households/parents: It refers to head of households with at least 13 years of education. The education of the head of household is also used to describe the household. In other words, households with head of household with at least 13 years of education are referred to as high school graduate households.
@@
Within household average years of education: It is the average education of children between 25 and 40 years for a given household. It serves as a metric for accessing the average quality of children within the household. A related variable is the Within household total years of education, which is the simple sum of children’s years of education. It is used as a proxy for the household’s total investment in education. The within household total and average years of education of children are denoted by $q_{T}$ and $\bar{q}$ respectively.

### v3

@@ Abstract
This paper examines how gender, birth order, and innate ability shape within-household disparities in children’s educational attainment in developing countries. Using data from Benin, I find that in households with non-educated parents, gender and birth order drive over two-thirds of the average educational attainment disparities among adult children, while their influence decreases to one-third in households with college-educated parents. Furthermore, average inequality, measured by the range of children’s educational attainment is twice as high among non-educated parents compared to college educated parents. I propose and estimate a structural model of educational attainment choices within-family. Using the model, I show that the absence of gender and birth order effects does not lead to a significant reduction in the average within-family disparities in children’s educational attainment. Additionally, in theory, ensuring that every child has at least one year of education lowers average within-family educational inequality. Yet, even in this scenario, daughters tend to receive less education than sons, and practical efforts to achieve universal entry are less effective than the theoretical model.
The latest version of this paper is available here.
@@ 1 Introduction
The trade-off between the number of children one chooses to have and the average education of those children has been a subject of extensive research. The quantity-quality trade-off model predicts a negative relationship between the quantity and the average quality of children. Numerous empirical studies have provided support for this model in various contexts, consistently revealing a negative relationship between the number of children, and the average educational attainment of those children (Becker and Lewis (1973), Montgomery (1995) and Li et al. (2008)). The current state of the literature on fertility choices (Conley and Glauber (2006), Maralani (2008), Li et al. (2008), Weng et al. (2019)) states that rich and educated families tend to have fewer children while allocating greater investments in the education of those children, in comparison to less affluent and less educated families. However, in contexts where fertility can not be controlled by parents— low access to contraceptive methods— or when the quality of and access to public education system is poor or there is no clear understanding of the economic benefits of schooling, this trade-off seems to vanish (Montgomery (1995), Black et al. (2005), Maralani (2008)).
@@
When we extend the analysis beyond the choice of the aggregate education of children to how it is distributed among children within a family, we evidence that, not only does the average change across households, but so does the variance of education. In addition, the variance is most likely non-zero for a majority of households when there are budget constraints and no or poorly enforced compulsory education laws. In such circumstances, the educational outcomes of children are strongly influenced by their observed and unobserved individual characteristics, leading to disparities in the amount of education they receive. Several studies have shed light on these disparities, with findings indicating that girls tend to receive less education due to factors such as gender bias or gender preference (as documented by Biswas (2000) and Ota and Moffatt (2007)). Additionally, birth order can play a role in the educational opportunities afforded to children, with elder siblings benefiting or facing disadvantages, as observed in studies like Ota and Moffatt (2007), Weng et al. (2019), Fergusson et al. (2006), De Haan (2010), Moshoeshoe et al. (2016), and Esposito et al. (2020). Furthermore, children with higher abilities are more likely to receive increased educational opportunities and educational attainment, as suggested by the research of Becker and Tomes (1976), Dizon-Ross (2019), and Giannola (2024). However, there exists a notable gap in the research landscape, as there are few studies that comprehensively analyze all these various sources of disparities within the same analytical framework. Such a framework, capable of simultaneously examining gender-based differences, birth order effects, and the impact of individual unobserved abilities on educational inequality within-households, holds the potential for the analysis of the effectiveness of educational counterfactual scenarios in reducing inequality. Specifically, it enables us to explore how these factors interact with one another to shape inequality. Furthermore, this comprehensive approach is crucial for estimating the proportion of inequality attributable to gender and birth order and the portion caused by other unobserved factors such as the cognitive ability of a child compared to their siblings. The former allows us to design educational policies which target inequality due to gender or birth order effects.
This micro-level analysis of educational disparities is particularly relevant to the broader context of educational challenges in Sub-Saharan Africa, where lack of education remains a critical obstacle to development. Despite numerous reforms, the region continues to struggle with high illiteracy rates, significant educational inequality, and have a substantial proportion of the world’s out-of-school children (UNESCO (2021)). While primary school completion rates are approaching or exceeding 90% in most regions globally, Sub-Saharan Africa lags behind, with only two out of three children completing primary school (UNESCO (2021)). Educational disparities are the results of various barriers, which can be categorized into three types: situational (life circumstances), dispositional (personal attitudes), and institutional (structural conditions) (UNESCO (2021)). A crucial but often overlooked factor contributing to these disparities is the within-family inequality in children’s educational attainment, which accounts for approximately 40% of the variation in educational attainment in the developing world (Giannola (2024)).
Using this regional evidence of educational inequality as a foundation, this paper focuses on the specific context of Benin. Benin is an ideal location for this study for three main reasons: the non enforcement of compulsory education laws, the significant variability in educational attainment among individuals within the same household, and the observed disparities in educational opportunities based on gender and birth order. For this exercise, I focused on households where there are two adult children residing in the same household as their parents [^3] The extension to households with more than 2 children is straightforward. All the estimates and analysis in the paper are also done after including households with 3 children and the results are presented in the appendix of the paper.. Among those households there are differences in how much educational resources parents have to distribute and the education attainment of the head of household. Taking that into account, I perform my analysis on households with different observed characteristics separately.
In the first part of the paper, I establish two key stylized facts about intra-household educational inequality. First, there is a non-linear, hump-shaped relationship between household-level mean and variance of educational attainment. Second, I decompose the average within-household educational attainment disparities into gender, birth order, and unobserved residual effects, finding that among households with non-educated heads and one child of each gender, over two-thirds of the average disparities in children’s educational attainment is due to gender and birth order, while among college-educated parents, only one-third is due to these factors. Furthermore, average inequality, measured by the range of children’s educational attainment is twice as high among non-educated parents compared to college educated parents. I then propose a structural model of household educational resources allocation to rationalize the observed inequality with budget constraints and parents’ preferences. The model accounts for gender effects by introducing a difference in the marginal utility of education between sons and daughters, with the marginal utility being higher for sons when birth order and other unobserved factors are held constant. Additionally, birth order effects are integrated through variations in the marginal cost of education by birth order, holding gender and unobserved factors fixed. Each child’s unique unobserved characteristics that influence their educational outcomes are modeled through an unobserved weight on the benefit derived from that child’s education. I estimate the model using the Simulated Method of Moments, and analyze diverse counterfactual scenarios in how they affect average inequality and the share of gender and birth order effects in the average inequality among non-educated parents. The first counterfactual (1) which theoretically set gender and birth order effects to zero; did not lead to a significant reduction in the average inequality in the sample. The second counterfactual (2) which theoretically removed barriers to school entry for all children reduces the average inequality in the sample by about half, but does not reduce the share of gender effect. The third counterfactual (3) which is more practical, relaxes the financial burden of non-educated parents to the same level as their college-educated counterparts. This only leads to a 10% reduction in the average inequality in the sample, with more than 50% of the remaining inequality due to gender effects.
@@
The subsequent sections of this paper are structured as follows. In Section 2, I present an overview of the data used for this study, while Section 3 presents key empirical evidence and stylized facts derived from the data description. Section 4 is dedicated for the model’s setup, outlining the estimation strategy for key parameters, and describing the inference and estimation procedures employed in this study. Lastly, in Section 5, I present counterfactual analysis to further explore the implications of my findings, and Section 7 concludes the paper.
Other relevant literature
@@
This paper contributes to the literature on within-household schooling decision, particularly factors influencing parents’ distribution decision of education resources among siblings. A key determinant of these distribution decisions is the gender of the child and the gender composition of the household. Previous research has shown that daughters are less likely to receive some education; or have lower educational attainment on average. Studies have shown that, while the presence of elder sisters tends to increase the likelihood of schooling, the presence of younger brothers may decrease it (Biswas (2000), Ota and Moffatt (2007), Ombati and Ombati (2012), Osadan and Burrage (2014), Psaki et al. (2018)). Another influential factor is the birth order of children, with mixed findings in previous studies. Some papers suggest a positive effect of birth order on children’s education (Ota and Moffatt (2007), Weng et al. (2019)), while others have shown that later-born children have lower educational attainment (Fergusson et al. (2006), De Haan (2010), Moshoeshoe et al. (2016), Esposito et al. (2020)). Finally, a child’s innate ability or talent plays a role in parental distribution decisions. Studies have demonstrated that parents invest more in the human capital of high-ability children and allocate more nonhuman capital to low-ability children (Becker and Tomes (1976), Dizon-Ross (2019), Giannola (2024)). When parents are compelled to invest in the nonhuman capital— for example inheritance in form of land or financial assets— of low-ability children, this leads to an inefficient equilibrium, where the investment in the human capital of high-ability children is not optimized (Nerlove et al. (1984)). This paper adds to this existing literature in two significant ways. First, it examines a context where parents are not constrained to compensate lower ability children by investing in their nonhuman capital but, instead, rely on family taxes (Wantchekon et al. (2015)). Second, this paper proposes a household educational resources distribution model which allows for a more flexible analysis of the distribution of education resources within the household. In this model, the assumption of equal distribution is relaxed, enabling a detailed exploration of the interactions between gender and, birth order effects, and the innate abilities of children in influencing household distribution decisions.
This paper also contributes to the literature on educational Kuznets curve theory ( Londoño (1990) and Ram (1990), Thomas et al. (2003), Morrisson and Murtin (2013)). Previous studies have analyzed the relationship between the mean and variance of education using cross country data or within-country time series data. This paper contributes to that literature by analyzing the relationship between the mean and variance of education using within-country cross household data. Specifically, it shows that in Benin at the household level; the relationship mean-variance of education is inverted U shaped with the peak occurring at approximately 7 years of education..
@@
Finally, this paper contributes to the literature on within-household inequality in children’s human capital (Giannola (2024)). Giannola (2024) has shown in the context of India that observed inequality within-households is partly explained by parents investing more in the human capital of high-achieving children, especially when they are financially constrained. This behavior stems from the fact that parents are not particularly averse to inequality and tend to reinforce the gap in learning created by innate ability rather than correcting it. This paper contributes to that literature by first building upon the result that parents unequally invest in the human capital of high-achieving children in contexts where the education system is better tailored to serve high-achieving students. Second, this paper interacts with that result and examines how it relates to other sources of inequality, such as gender and birth order.
@@ 2.1 Sample and Data
In this section, I present the data used. I use data from the 2013 Population and Habitation Census of Benin. This census data provides information on all households and their members. It is conducted by the National Bureau of Statistics of Benin, and has information at both the household and individual levels. For the purpose of this paper, the focus is directed towards individuals who identify themselves as the children of household heads, enabling to get information on parental  [^4] Parents here refers to one of the parents, either the mother or the father. This because it is not possible to have both for household with single parents and to identify the biological mother for polygamous households. and sibling characteristics for a sub-sample of siblings. The variable “Number of children” represents the observed number of children within each household  [^5] It does not include children who moved out of the family house before the census.. For the primary analysis, only households with children aged between 25 and 40 years are included. This age range is chosen to ensure that the children have either completed their education or nearly achieved maximum educational attainment. The summary statistics of all individuals between 25 and 40 years living in the same household as their parents is presented in Table 9 in Appendix A.
The inequality analysis focuses on households with at least 2 such children falling within the specified age range and at least one child with some educational attainment. This specific condition on the sample is motivated by the goal to examine the reasons for providing equal education to all children, as opposed to the alternative of not educating any children. In particular households with only non-educated children do not offer any information about the distribution of education resources which this paper aims to analyze. The resulting sample comprises approximately $90,000$ individuals and $\approx 33,000$ households, serving as the basis for further investigation. The sample description is as follows:
  1. 1.
    Sample 0: All households with children between 25-40 yrs old ($\approx 160,000$ households)
  2. 2.
    Sample 1: Households in sample 0 with at least 2 children between 25 and 40 years old. ($\approx$ 51,600 households)
  3. 3.
    Sample 2: Households in sample 1 with at least one educated child between 25 and 40 years old. ($\approx$ 32,800 households)
I use sample 1 for stylized facts, estimation of the model, counterfactual analysis, and comparative statics, and sample 2 only for counterfactual analysis, and comparative statics.
@@ 2.2 Key variables and measurement
The data set contains several key variables used in this paper, including gender, age, religion, area of residence, family size, household wealth index, and educational attainment of individuals, as well as their parents’ and a subset of their siblings’ variables. Apart from these variables, I also created measures for within-household inequality, within-household average years of educational attainment of children, and gender composition of children within a household. A description of each variable and their measurement is as follows:
Within household inequality: Is the disparity in the educational attainment of children within a given household. It is measured by the within-household range of children’s education attainment for households with 2 children. For households with more than 2 children, it is measured by the standard deviation of education.
@@
Within-household average years of educational attainment of children: Is the average education of children between 25 and 40 years for a given household. It serves as a metric for accessing the average quality of children within the household, and is used as proxy for educational resources available. A related variable is the Within-household total years of educational attainment of children, which is the simple sum of children’s years of education. It is used as a proxy for the household’s total investment in education. The within-household total and average years of educational attainment of children are denoted by $q_{T}$ and $\bar{q}$ respectively.